# Accelerated Sinkhorn Algorithms for Partial Optimal Transport

### v1

@@ Abstract
Partial Optimal Transport (POT) addresses the problem of transporting only a fraction of the total mass between two distributions, making it suitable when marginals have unequal size or contain outliers. While Sinkhorn-based methods are widely used, their complexity bounds for POT remain suboptimal and can limit scalability. We introduce Accelerated Sinkhorn for POT (ASPOT), which integrates alternating minimization with Nesterov-style acceleration in the POT setting, yielding a complexity of $\mathcal{O}(n^{7/3}\varepsilon^{-5/3})$. We also show that an informed choice of the entropic parameter $\gamma$ improves rates for the classical Sinkhorn method. Experiments on real-world applications validate our theories and demonstrate the favorable performance of our proposed methods.
@@ 1 Introduction
Optimal Transport (OT) has been a core tool to compare probability distributions, with broad impact from economics to machine learning [11, 12, 18, 23]. Its success stems from efficient computational methods [3, 10, 11, 20] and the role of entropic regularization, which improves statistical properties and enables seamless integration into deep learning frameworks [2, 10, 12, 21, 5, 7, 31]. Classical OT, however, requires equal mass and is sensitive to outliers [29]. POT addresses these issues by relaxing mass constraints, making it more robust in practical tasks such as image analysis and distributional learning [35, 36, 34, 17, 22, 6]. However, existing computational approaches remain unsatisfactory. A practical path to POT is to reduce it to a balanced OT instance by adding dummy nodes and then apply Sinkhorn‐type iterations [6]. While straightforward, naive variants can be infeasible [17]. A feasible formulation with guarantees was established in [22], yielding a Sinkhorn complexity of $\mathcal{O}\left(n^{2}/\varepsilon^{4}\right)$ and proposing APDAGD with $\mathcal{O}\left(n^{2.5}/\varepsilon\right)$. However, applied researchers note that tuning the Sinkhorn parameter often yields stronger performance, while APDAGD’s extra $\mathcal{O}(n^{0.5})$ factor limits its scalability. Thus, our central questions are:
@@ 3.3 Theoritical lemmas and theorems
We refer the reader to Section 8 for the construction of the matrix $A$ (via slack variables turning POT into a linear programm), complete proofs, extra algorithms, and further experiments. From [22], we recall the bound on the dual variables $z^{\star}$:
@@ 4.1 Color transfer
Experimental setup. We follow the setup of [4]. AI-generated images (Fig. 1) are represented in $\mathbb{R}^{3}$ by RGB values of pixels and quantized to $n=800$ colors using $k$ means, producing histograms $a,b$ as source and target distributions. Both are normalized by $\max\{\|a\|_{1},\|b\|_{1}\}$ so total mass $\leq 1$. The cost is $C_{ij}=\|a_{i}-b_{j}\|^{2}/\max(C)$, and the transport budget is $s=0.2\cdot\min(\sum a,\sum b)$. We compare ASPOT, feasible Sinkhorn, and APDAGD, each run for 1500 iterations with $\gamma=10^{-3}$ and $\texttt{tol}=10^{-7}$. After computing a plan $X$, barycentric projection $\hat{a}_{i}=\sum_{j}X_{ij}b_{j}/\sum_{j}X_{ij}$ recolors pixels assigned to centroid $a_{i}$, transforming the new image.
@@
ASPOT results. Fig. 1 compares the three methods on color transfer. The convergence curves show that ASPOT rapidly reduces transport cost, reaching good solutions far earlier than the others. Sinkhorn progresses steadily but much slower, while APDAGD stalls for many iterations before improving. As a result, when ASPOT is already near optimal, the other methods still lag by a gap consistent with ASPOT’s better $\varepsilon$-dependence. The visual results confirm that APDAGD and Sinkhorn yield almost identical outputs, both washed out, lacking detail, with incomplete recoloring and visible artifacts, reflecting their slow convergence. By contrast, ASPOT transfers cool tones more clearly and preserves source structure, producing sharper and more natural images. Overall, ASPOT both converges faster and produces better visual quality than the baselines.
@@ 6 Acknowledgement
The authors express their profound gratitude to Mr. Quang Minh Nguyen (MIT, nmquang@mit.edu) and Mr. Hoang Huy Nguyen (Georgia Tech, hnguyen455@gatech.edu) for their invaluable input in this project.

### v2

@@ Abstract
Partial Optimal Transport (POT) addresses the problem of transporting only a fraction of the total mass between two distributions, making it suitable when marginals have unequal size or contain outliers. While Sinkhorn-based methods are widely used, their complexity bounds for POT remain suboptimal and can limit scalability. We introduce Accelerated Sinkhorn for POT (ASPOT), which integrates alternating minimization with Nesterov-style acceleration in the POT setting, yielding a complexity of $\mathcal{O}(n^{7/3}\varepsilon^{-5/3})$. We also show that an informed choice of the entropic parameter $\gamma$ improves rates for the classical Sinkhorn method. Experiments on real-world applications validate our theories and demonstrate the favorable performance of our proposed method. Our code is open-source at the link github.
@@ 1 Introduction
Optimal Transport (OT) has been a core tool to compare probability distributions, with broad impact from economics to machine learning [11, 12, 18, 23]. Its success stems from efficient computational methods [3, 10, 11, 20] and the role of entropic regularization, which improves statistical properties and enables seamless integration into deep learning frameworks [2, 10, 12, 21, 5, 7]. Classical OT, however, requires equal mass and is sensitive to outliers [29]. POT addresses these issues by relaxing mass constraints, making it more robust in practical tasks such as image analysis, few-shot learning and distributional learning [35, 36, 34, 17, 22, 6]. However, existing computational approaches remain unsatisfactory. A practical path to POT is to reduce it to a balanced OT instance by adding dummy nodes and then apply Sinkhorn‐type iterations [6]. While straightforward, naive variants can be infeasible [17]. A feasible formulation with guarantees was established in [22], yielding a Sinkhorn complexity of $\mathcal{O}\left(n^{2}/\varepsilon^{4}\right)$ and proposing APDAGD with $\mathcal{O}\left(n^{2.5}/\varepsilon\right)$. However, applied researchers note that tuning the Sinkhorn parameter often yields stronger performance, while APDAGD’s extra $\mathcal{O}(n^{0.5})$ factor limits its scalability. Thus, our central questions are:
@@ 3.3 Theoritical lemmas and theorems
We refer the reader to [[30], Appendix] for the construction of the matrix $A$ (via slack variables turning POT into a linear programm), complete proofs, extra algorithms, and further experiments. From [22], we recall the bound on the dual variables $z^{\star}$:
@@ 4.1 Color transfer
Experimental setup. We follow the setup of [4]. AI-generated images (Fig. 1) are represented in $\mathbb{R}^{3}$ by RGB values of pixels and quantized to $n=800$ colors using $k$ means, producing histograms $a,b$ as source and target distributions. Both are normalized by $\max\{\|a\|_{1},\|b\|_{1}\}$ so total mass $\leq 1$. The cost is $C_{ij}=\|a_{i}-b_{j}\|^{2}/\max(C)$, and the transport budget is $s=0.2\cdot\min(\sum a,\sum b)$. We compare ASPOT, feasible Sinkhorn, and APDAGD, each run for 1500 iterations with $\gamma=10^{-3}$ and $\texttt{tol}=10^{-7}$. After computing a plan $X$, barycentric projection $\hat{a}_{i}=\sum_{j}X_{ij}b_{j}/\sum_{j}X_{ij}$ recolors pixels assigned to centroid $a_{i}$, transforming the new image. ASPOT results. Fig. 1 compares the three methods on color transfer. The convergence curves show that ASPOT rapidly reduces transport cost, reaching good solutions far earlier than the others. Sinkhorn progresses steadily but much slower, while APDAGD stalls for many iterations before improving. As a result, when ASPOT is already near optimal, the other methods still lag by a gap consistent with ASPOT’s better $\varepsilon$-dependence. The visual results confirm that APDAGD and Sinkhorn yield almost identical outputs, both washed out, lacking detail, with incomplete recoloring and visible artifacts, reflecting their slow convergence. By contrast, ASPOT transfers cool tones more clearly and preserves source structure, producing sharper and more natural images. Overall, ASPOT both converges faster and produces better visual quality than the baselines.
@@ 6 Acknowledgement
The authors express their profound gratitude to Mr. Quang Minh Nguyen (MIT, nmquang@mit.edu) and Mr. Hoang Huy Nguyen (Georgia Tech, hnguyen455@gatech.edu) for their invaluable input in this project. This research is supported by VNUHCM-University of Information Technology’s Scientific Research Support Fund.